# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v3

@@ 3 Methods
Salt-and-pepper Noise The salt-and-pepper noise is a common multiplicative noise for images. The image can exhibit unnatural changes, such as black pixels in bright areas or white pixels in dark areas, specifically as a result of the signal disruption caused by sudden strong interference or bit transmission errors. In the Salt-and-pepper noise case, the entropy change is:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})\\ =&-\sum_{\boldsymbol{Z}\in\mathcal{Z}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Z},\boldsymbol{Y})\log p(\boldsymbol{Z},\boldsymbol{Y})+\sum_{\boldsymbol{Z}\in\mathcal{Z}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}\sum_{\boldsymbol{\epsilon}\in\mathcal{E}}p(\boldsymbol{Z\epsilon},\boldsymbol{Y})\log p(\boldsymbol{Z\epsilon},\boldsymbol{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z\epsilon},\boldsymbol{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{Z},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-H(\boldsymbol{\epsilon})\end{split}$ (20)
Obviously, the entropy change is smaller than 0, which indicates the complexity is increasing when injecting salt-and-pepper noise into the deep model. As can be foreseen, the salt-and-pepper noise is pure detrimental noise. More details and Salt-and-pepper added to the image level are in the supplementary.
@@ 4 Experiments
In this section, we conduct extensive experiments to explore the influence of various types of noises on deep learning models. We employ popular deep learning architectures, including both CNNs and ViTs, and show that the two kinds of deep models can benefit from the positive noise. We employ deep learning models of various scales, including ViT-Tiny (ViT-T), ViT-Small (ViT-S), ViT-Base (ViT-B), and ViT-Large (ViT-L) for Vision Transformers (ViTs), and ResNet-18, ResNet-34, ResNet-50, and ResNet-101 for ResNet architecture. The details of deep models are presented in the supplementary. Without specific instructions, the noise is injected at the last layer of the deep models. Note that for ResNet models, the number of macro layers is 4, and for each macro layer, different scale ResNet models have different micro sublayers. For example, for ResNet-18, the number of macro layers is 4, and for each macro layer, the number of micro sublayers is 2. The noise is injected at the last micro sublayer of the last macro layer for ResNet models. More experimental settings for ResNet and ViT are detailed in the supplementary.
@@ 4.1 Noise Setting
For linear transform noise, many possible quality $Q$ matrices could satisfy these constraints, forming a design space. Here, we adopt a simple concrete construction of $Q$ that we call a circular shift as a working example. In this construction, each original $Z_{i}$ is perturbed slightly by its immediate next neighbor $Z_{i+1}$. We can formally express the circular shift noise injection strategy as follows: Let the scalar hyperparameter $\alpha\in[0,1]$ define the perturbation strength. The quality matrix $Q$ is implemented as $Q=\alpha*U-\alpha*I$, where $U_{i,j}=\delta_{i+1,j}$ with $\delta_{i+1,j}$ representing the Kronecker delta indicator [18], and employing wrap-around (or “circular”) indexing. The concrete formation of the quality matrix in circular shift form is then formulated as:
@@
The parameter $\alpha$ represents the linear transform strength.
@@
$\begin{split}\begin{cases}max(X)&\text{ if }p<\alpha/2\\ min(X)&\text{ if }p>1-\alpha/2\end{cases}\end{split}$ (28)
where $p$ is a probability generated by a random seed, $\alpha\in[0,1)$, and $X$ is the representation of an image.
Table 1: ResNet with different kinds of noise on ImageNet. Vanilla means the vanilla model without noise. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket.
@@
Table 2: ViT with different kinds of noise on ImageNet. Vanilla means the vanilla model without injecting noise. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket. Note ViT-L is overfitting on ImageNet [15] [54].
@@
Table 3: Comparison between Positive Noise Empowered ViT with other ViT variants. Top 1 Accuracy is shown in percentage. Here PN is the positive noise, i.e., linear transform noise.
@@
ViT-B [15] 84.33 86M 224 $\times$ 224 ImageNet 21k
DeiT-B [59] 85.70 86M 224 $\times$ 224 ImageNet 21k
SwinTransformer-B [33] 86.40 88M 384 $\times$ 384 ImageNet 21k
DaViT-B [14] 86.90 88M 384 $\times$ 384 ImageNet 21k
MaxViT-B [60] 88.82 119M 512 $\times$ 512 JFT-300M (Private)
ViT-22B [12] 89.51 21743M 224 $\times$ 224 JFT-4B (Private)
NoisyViT-B (ViT+B+PN) 89.99 86M 224 $\times$ 224 ImageNet 21k
NoisyViT-B (ViT+B+PN) 91.37 86M 384 $\times$ 384 ImageNet 21k
Table 4: Top 1 accuracy on ImageNet V2 with positive linear transform noise.
Model Top1 Acc. Params. Image Res. Pretrained Dataset
NoisyViT-B 82.20 86M 224 $\times$ 224 ImageNet 21k
NoisyViT-B 84.80 86M 384 $\times$ 384 ImageNet 21k
@@ 4.2 Image Classification Results
We implement extensive experiments on large-scale datasets such as ImageNet [13] and small-scale datasets such as TinyImageNet [27] using ResNets and ViTs.
@@ 4.2.1 CNN Family
The results of ResNets with different noises on ImageNet are in Table 1. As shown in the table, with the design of linear transform noise to be positive noise (PN), ResNet improves the classification accuracy by a large margin. While the salt-and-pepper, which is theoretically harmful noise (HN), degrades the models. Note we did not utilize data augmentation techniques for ResNet experiments except for normalization. The significant results show that positive noise can effectively improve classification accuracy by reducing task complexity.
@@ 4.2.2 ViT Family
The results of ViT with different noises on ImageNet are in Table 2. Since the ViT-L is overfitting on the ImageNet [15] [54], the positive noise did not work well on the ViT-L. As shown in the table, the existence of positive noise improves the classification accuracy of ViT by a large margin compared to vanilla ViT. The comparisons with previously published works, such as DeiT [59], SwinTransformer [33], DaViT [14], and MaxViT [60], are shown in Table 3, and our positive noise-empowered ViT achieved the new state-of-the-art result. Note that the JFT-300M and JFT-4B datasets are private and not publicly available [56], and we believe that ViT large and above will benefit from positive noise significantly if trained on larger JFT-300M or JFT-4B, which is theoretically supported in section 4.4.
@@ 4.3 Ablation Study
We also proactively inject noise into variants of ViT, such as DeiT [59], Swin Transformer [33], BEiT [3], and ConViT [11], and the results show that positive noise could benefit various variants of ViT by improving classification accuracy significantly. The results of injecting noise to variants of ViT are reported in the supplementary. We also did ablation studies on the strength of linear transform noise and the injected layer. The results are shown in Fig. 2. We can observe that the deeper layer the positive noise injects, the better prediction performance the model can obtain. There are reasons behind this phenomenon. First, the latent features of input in the deeper layer have better representations than those in shallow layers; second, injection to shallow layers reduces less task complexity because of trendy replacing Equation 10 with Equation 9. More results on the small dataset TinyImageNet can be found in the supplementary.
Additionally, we tested the positive linear transformation noise on another popular dataset, the ImageNet V2. The corresponding results are reported in Table 4.
@@
NoisyViT-B+Optimal Q 95.65 86M 384 $\times$ 384 ImageNet 21k
Table 6: Comparison with various ViT-based methods on Office-Home.
@@
ViT-B [15] 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-B [67] 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-B [66] 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B [57] 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
@@
4.4 Optimal Quality Matrix
$Q_{optimal}=\mathrm{diag}\left(\frac{1}{k+1}-1,\dots,\frac{1}{k+1}-1\right)+\frac{1}{k+1}\boldsymbol{1}_{k\times k}$ (29)
$\bigtriangleup S(\mathcal{T},Q_{optimal}\boldsymbol{X})=(k-1)\log{(k+1)}$ (30)
[…]
4.5 Domain Adaption Results
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions [42] [63]. Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home [62] and VisDA2017 [43]. Detailed datasets introduction and experiments training settings are in the supplementary. The objective function is borrowed from TVT [67], which is the first work that adopts Transformer-based architecture for UDA. The results are shown in Table 6 and 7. The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results demonstrate that positive noise also works in domain adaptation tasks, where out-of-distribution (OOD) data exists.
@@ 5 Conclusion
This study delves into the influence of entropy change on learning systems, achieved by proactively introducing various types of noise into deep models. Our work conducts a comprehensive investigation into the impact of common noise types, such as Gaussian noise, linear transform noise, and salt-and-pepper noise, on deep learning models. Notably, we demonstrate that, under specific conditions, linear transform noise can positively affect deep models. The experimental results show that injecting positive noise into the latent space significantly enhances the prediction performance of deep models in image classification tasks, leading to new state-of-the-art results on ImageNet. These findings hold broad implications for future research and the potential development of more accurate models for improved real-world applications.
6 Potential Impact
The proposed theory exhibits versatility, opening avenues for exploring the application of positive noise across diverse deep learning tasks in computer vision and natural language processing. While our current findings and theoretical analysis are centered around addressing classification problems in the computer vision domain, their implications extend far beyond. Specifically, the insights hold substantial promise for various learning tasks, including those involving large language models (LLMs). The optimal quality matrix derived from the linear transform noise suggests a potential for more effective model enhancement with larger datasets. This finding implies that the principles uncovered in our study may contribute to refining the strategies for improving large-scale language models. The adaptability of the proposed theory to both computer vision and NLP tasks marks a promising avenue for future research. A potential concern is that those who possess high-quality large-scale datasets may primarily benefit from our research.
@@ Appendix A Theoretical Foundations of Task Entropy
This section provides the theoretical foundations of task entropy, quantifying the complexity of learning tasks. The concept of task entropy was first proposed for the image level and formulated as [30]:
$H(\mathcal{T};\boldsymbol{X})=-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X})\log{p(\boldsymbol{Y}|\boldsymbol{X})}$ (31)
The image $\boldsymbol{X}$ in the dataset are supposed to be independent of each other, as are the labels $\boldsymbol{Y}$. However, $\boldsymbol{X}$ and $\boldsymbol{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$. Essentially, the task entropy is the entropy of $p(Y|X)$. Following the principle of task entropy, compelling evidence suggests that diminishing task complexity via reducing information entropy can enhance overall model performance [30, 69].
@@
$H(\mathcal{T};\boldsymbol{Z})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})$ (32)
where $\boldsymbol{Z}$ are the embeddings of the images $\boldsymbol{X}$. Here, we assume that the embedding $\boldsymbol{Z}$ and the vectorized label $\boldsymbol{Y}$ follow a multivariate normal distribution. We can transform the unknown distributions of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ to approximately conform to normality by utilizing techniques such as Box-Cox transformation, log transform, etc., as described in [5] and [17] if they are not initially normal. After approximate transformation, the distribution of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ can be expressed as:
$\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\mu_{Z}},\boldsymbol{\Sigma_{Z}}),\boldsymbol{Y}\sim\mathcal{N}(\boldsymbol{\mu_{Y}},\boldsymbol{\Sigma_{Y}})$ (33)
@@
$\begin{split}&\boldsymbol{{\mu}_{Z}}=\mathbb{E}[\boldsymbol{Z}]=(\mathbb{E}[Z_{1}],\mathbb{E}[Z_{2}],...,\mathbb{E}[Z_{k}]])^{T}\\ &\boldsymbol{{\mu}_{Y}}=\mathbb{E}[\boldsymbol{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\boldsymbol{\Sigma_{Z}}=\mathbb{E}[(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})^{T}]\\ &\boldsymbol{\Sigma_{Y}}=\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\end{split}$ (34)
@@
Then the conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{Z}$ is also normally distributed [38] [24], which can be formulated as:
$\boldsymbol{Y}|\boldsymbol{Z}\sim\mathcal{N}(\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}={Z}}),var(\boldsymbol{Y}|\boldsymbol{Z}={Z}))$ (35)
where $\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}=Z})$ is the mean of the label set $\boldsymbol{Y}$ given a sample $\boldsymbol{Z}=Z$ from the embeddings, and $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ is the variance of $\boldsymbol{Y}$ given a sample from the embeddings. The conditional mean $\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]$ and conditional variance $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ can be calculated as:
$\boldsymbol{\mu}_{\boldsymbol{Y}|\boldsymbol{Z}=Z}=\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]=\boldsymbol{{\mu}_{Y}}+\boldsymbol{\Sigma_{YZ}}\boldsymbol{{\Sigma}^{-1}_{Z}}(Z-\boldsymbol{\mu_{Z}})$ (36)
$\boldsymbol{\Sigma}_{{\boldsymbol{Y|Z}}=Z}=var(\boldsymbol{Y}|\boldsymbol{Z}=Z)=\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}$ (37)
where $\boldsymbol{\Sigma_{YZ}}$ and $\boldsymbol{\Sigma_{ZY}}$ are the cross-covariance matrices between $\boldsymbol{Y}$ and $\boldsymbol{Z}$, and between $\boldsymbol{Z}$ and $\boldsymbol{Y}$, respectively, and $\boldsymbol{\Sigma^{-1}_{Z}}$ denotes the inverse of the covariance matrix of $\boldsymbol{Z}$.

### v4

@@ 3 Methods
Salt-and-pepper Noise. The salt-and-pepper noise is a common multiplicative noise for images. The image can exhibit unnatural changes, such as black pixels in bright areas or white pixels in dark areas, specifically as a result of the signal disruption caused by sudden strong interference or bit transmission errors. In the Salt-and-pepper noise case, the entropy change is:
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&H(\bm{Y};\bm{Z})-H(\bm{Z})-(H(\bm{Y};\bm{Z}\bm{\epsilon})-H(\bm{Z}))\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}\bm{\epsilon})\\ =&-\sum_{\bm{Z}\in\mathcal{Z}}\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Z},\bm{Y})\log p(\bm{Z},\bm{Y})+\sum_{\bm{Z}\in\mathcal{Z}}\sum_{\bm{Y}\in\mathcal{Y}}\sum_{\bm{\epsilon}\in\mathcal{E}}p(\bm{Z\epsilon},\bm{Y})\log p(\bm{Z\epsilon},\bm{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{Z\epsilon},\bm{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{Z},\bm{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\bm{\epsilon})}\right]\\ =&-H(\bm{\epsilon})\end{split}$ (16)
Obviously, the entropy change is smaller than 0, which indicates the complexity is increasing when injecting salt-and-pepper noise into the deep model. As can be foreseen, the salt-and-pepper noise is pure detrimental noise. More details and Salt-and-pepper added to the image level are in App. D.
@@ 4 Experiments
In this section, we conduct extensive experiments to explore the influence of various types of noises on deep learning models. We employ popular deep learning architectures, including both CNNs and ViTs, and show that the two kinds of deep models can benefit from the positive noise. We employ deep learning models of various scales, including ViT-Tiny (ViT-T), ViT-Small (ViT-S), ViT-Base (ViT-B), and ViT-Large (ViT-L) for Vision Transformers (ViTs), and ResNet-18, ResNet-34, ResNet-50, and ResNet-101 for ResNet architecture. The details of deep models are presented in the supplementary. Without specific instructions, the noise is injected at the last layer of the deep models. Note that for ResNet models, the number of macro layers is 4, and for each macro layer, different scale ResNet models have different micro sublayers. For example, for ResNet-18, the number of macro layers is 4, and for each macro layer, the number of micro sublayers is 2. The noise is injected at the last micro sublayer of the last macro layer for ResNet models. More experimental settings for ResNet and ViT are detailed in the supplementary. While this work primarily focuses on image classification and domain adaptation, we additionally explored other related tasks: Domain Generalization (App F.8) and Text Classification (App F.9) to assess broader applicability of NoisyNN.
@@ 4.1 Noise Setting
For linear transform noise, many possible quality $Q$ matrices could satisfy these constraints, forming a design space. Here, we adopt a simple concrete construction of $Q$ that we call a circular shift as a working example. In this construction, each original $Z_{i}$ is perturbed slightly by its immediate next neighbor $Z_{i+1}$. We can formally express the circular shift noise injection strategy as follows: Let the scalar hyperparameter $\alpha\in[0,1]$ define the perturbation strength. The quality matrix $Q$ is implemented as $Q=\alpha*U-\alpha*I$, where $U_{i,j}=\delta_{i+1,j}$ with $\delta_{i+1,j}$ representing the Kronecker delta indicator Frankel (2011), and employing wrap-around (or “circular”) indexing. The concrete formation of the quality matrix in circular shift form is then formulated as:
@@
where the parameter $\alpha$ represents the linear transform strength.
@@
$\begin{split}\begin{cases}max(Z)&\text{ if }p<\alpha/2\\ min(Z)&\text{ if }p>1-\alpha/2\end{cases}\end{split}$ (19)
where $p$ is a probability generated by a random seed, $\alpha\in[0,1)$, and $Z$ is the latent representation of an image.
Table 1: ResNet with different kinds of noise on ImageNet. Vanilla means the vanilla model without noise injection. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket.
@@
Table 2: ViT with different kinds of noise on ImageNet. Vanilla means the vanilla model without injecting noise. Accuracy is shown in percentage. Gaussian noise used here is subjected to standard normal distribution. Linear transform noise used in this table is designed to be positive noise. The difference is shown in the bracket.
@@
Table 3: Comparison between Positive Noise Empowered ViT with other ViT variants. Top 1 Accuracy is shown in percentage. Here, PN refers to positive noise, specifically a linear transformation designed to be positive noise.
@@
ViT-B Dosovitskiy et al. (2020) 84.33 86M 224 $\times$ 224 ImageNet 21k
DeiT-B Touvron et al. (2021) 85.70 86M 224 $\times$ 224 ImageNet 21k
SwinTransformer-B Liu et al. (2021) 86.40 88M 384 $\times$ 384 ImageNet 21k
DaViT-B Ding et al. (2022) 86.90 88M 384 $\times$ 384 ImageNet 21k
MaxViT-B Tu et al. (2022) 88.82 119M 512 $\times$ 512 JFT-300M (Private)
ViT-22B Dehghani et al. (2023) 89.51 21743M 224 $\times$ 224 JFT-4B (Private)
NoisyViT-B (ViT-B+PN) 89.99 86M 224 $\times$ 224 ImageNet 21k
NoisyViT-B (ViT-B+PN) 91.37 348M 384 $\times$ 384 ImageNet 21k
@@ 4.2 Image Classification Results
We implement extensive experiments on large-scale datasets such as ImageNet Deng et al. (2009) and small-scale datasets such as TinyImageNet Le & Yang (2015a) using ResNets and ViTs.
@@ 4.2.1 CNN Family
The results of ResNets with different noises on ImageNet are in Table 1. As shown in the table, with the design of linear transform noise to be positive noise (PN), ResNet improves the classification accuracy by a large margin. While the salt-and-pepper, which is theoretically harmful noise (HN), degrades the models. Note we did not utilize data augmentation techniques for ResNet experiments. The significant results show that positive noise can effectively improve classification accuracy by reducing task complexity.
@@ 4.2.2 ViT Family
The results of ViT with different noises on ImageNet are in Table 2. Since the ViT-L is overfitting on the ImageNet Dosovitskiy et al. (2020) Steiner et al. (2021), the positive noise did not work well on the ViT-L. As shown in the table, the existence of positive noise improves the classification accuracy of ViT by a large margin compared to vanilla ViT. The comparisons with previously published works, such as DeiT Touvron et al. (2021), SwinTransformer Liu et al. (2021), DaViT Ding et al. (2022), and MaxViT Tu et al. (2022), are shown in Table 3, and our positive noise-empowered ViT achieved the new state-of-the-art result. Note that the JFT-300M and JFT-4B datasets are private and not publicly available Sun et al. (2017), and we believe that ViT large and above will benefit from positive noise significantly if trained on larger datasets, which is theoretically supported in section 4.4.
@@ 4.3 Ablation Study
We also proactively inject noise into variants of ViT, such as DeiT Touvron et al. (2021), Swin Transformer Liu et al. (2021), BEiT Bao et al. (2021), and ConViT d’Ascoli et al. (2021), and the results show that positive noise could benefit various variants of ViT by improving classification accuracy significantly. The results of injecting noise to variants of ViT are reported in the supplementary Table 10. We also conducted ablation studies on the strength of linear transform noise and the injected layer. The results are shown in Fig. 2. We can observe that the deeper layer the positive noise injects, the better prediction performance the model can obtain. There are reasons behind this phenomenon. First, the latent features of input in the deeper layer have more abstract representations than those in shallow layers; second, injection to shallow layers reduces less task complexity because of trendy replacing Equation 8 with Equation 7. More results on the small dataset TinyImageNet can be found in the supplementary.
Additionally, we tested the positive linear transformation noise on another popular dataset, the ImageNet V2 Recht et al. (2019). The corresponding results are reported in Table 4.
Table 4: Top 1 accuracy on ImageNet V2 with positive linear transform noise that is designed to be positive noise.
Model Top1 Acc. Params. Image Res.
ViT-B 73.9 86M 224 $\times$ 224
NoisyViT-B 82.2 86M 224 $\times$ 224
NoisyViT-B 84.8 348M 384 $\times$ 384
4.4 Optimal Quality Matrix
$Q_{optimal}=\mathrm{diag}\left(\frac{1}{k+1}-1,\dots,\frac{1}{k+1}-1\right)+\frac{1}{k+1}\bm{1}_{k\times k}$ (20)
$\bigtriangleup S(\mathcal{T},Q_{optimal}\bm{X})=(k-1)\log{(k+1)}$ (21)
[…]
@@ 4.4 Optimal Quality Matrix
NoisyViT-B+Optimal Q 95.65 348M 384 $\times$ 384 ImageNet 21k
4.5 Domain Adaption Results
Table 6: Comparison with various ViT-based methods, i.e. ViT-B Dosovitskiy et al. (2020), TVT-B Yang et al. (2023a), CDTrans-B Xu et al. (2022) and SSRT-B Sun et al. (2022) on Office-Home.
@@ 4.5 Domain Adaption Results
ViT-B Dosovitskiy et al. (2020) 97.7 48.1 86.6 61.6 78.1 63.4 94.7 10.3 87.7 47.7 94.4 35.5 67.1
TVT-B Yang et al. (2023a) 92.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.9
CDTrans-B Xu et al. (2022) 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
SSRT-B Sun et al. (2022) 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
@@
Unsupervised domain adaptation (UDA) aims to learn transferable knowledge across the source and target domains with different distributions Pan & Yang (2009) Wei et al. (2018). Recently, transformer-based methods achieved SOTA results on UDA, therefore, we evaluate the ViT-B with the positive noise on widely used UDA benchmarks. Here the positive noise is the linear transform noise identical to that used in the classification task. The positive noise is injected into the last layer of the model, the same as the classification task. The datasets include Office Home Venkateswara et al. (2017) and VisDA2017 Peng et al. (2017). Detailed datasets introduction and experiments training settings are in the supplementary. The objective function is borrowed from TVT Yang et al. (2023a). The results are shown in Table 6 and 7 (the complete tables can be found in Tab. 21 and Tab. 22 in the Appendix). The NoisyTVT-B, i.e., TVT-B with positive noise, achieves better performance than existing works. These results demonstrate that positive noise also works in domain adaptation tasks, where out-of-distribution (OOD) data exists.
@@ 5 Conclusion
This study delves into the influence of entropy change on learning tasks, achieved by proactively introducing various types of noise into deep models. Our work conducts a comprehensive investigation into the impact of common noise types, such as Gaussian noise, linear transform noise, and salt-and-pepper noise, on deep learning models. Notably, we demonstrate that, under specific conditions, linear transform noise can positively affect deep models. The experimental results show that injecting positive noise into the latent space significantly enhances the prediction performance of deep models in image classification tasks, leading to new state-of-the-art results on ImageNet. These findings hold broad implications for future research and the potential development of more accurate models for improved real-world applications.
subsectionPotential Impact The proposed theory exhibits versatility, opening avenues for exploring the application of positive noise across diverse deep learning tasks in computer vision and natural language processing. While our current findings and theoretical analysis are centered around addressing problems in the image classification tasks in computer vision, their implications extend far beyond. The optimal quality matrix derived from the linear transform noise suggests a potential for more effective model enhancement with larger datasets. This finding implies that the principles uncovered in our study may contribute to refining the strategies for improving large-scale language models trained on large-scale datasets. A potential concern is that those who possess high-quality large-scale datasets may primarily benefit from our research.
@@ Appendix A Theoretical Foundations of Task Entropy
This section provides the theoretical foundations of task entropy, quantifying the complexity of learning tasks. The concept of task entropy was first proposed for the image level and formulated as Li (2022):
$H(\mathcal{T};\bm{X})=-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X})\log{p(\bm{Y}|\bm{X})}$ (22)
The image $\bm{X}$ in the dataset are supposed to be independent of each other, as are the labels $\bm{Y}$. However, $\bm{X}$ and $\bm{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$. Essentially, the task entropy is the entropy of $p(Y|X)$. Following the principle of task entropy, compelling evidence suggests that diminishing task complexity via reducing information entropy can enhance overall model performance Li (2022); Zhang et al. (2023).
@@
$H(\mathcal{T};\bm{Z})\coloneqq H(\bm{Y},\bm{Z})-H(\bm{Z})$ (23)
where $\bm{Z}$ are the embeddings of the images $\bm{X}$. Here, we assume that the embedding $\bm{Z}$ and the vectorized label $\bm{Y}$ follow a multivariate normal distribution. We can transform the unknown distributions of $\bm{Z}$ and $\bm{Y}$ to approximately conform to normality by utilizing techniques such as Box-Cox transformation, log transform, etc., as described in Box & Cox (1964) and Feng et al. (2014) if they are not initially normal. After approximate transformation, the distribution of $\bm{Z}$ and $\bm{Y}$ can be expressed as:
$\bm{Z}\sim\mathcal{N}(\bm{\mu_{Z}},\bm{\Sigma_{Z}}),\bm{Y}\sim\mathcal{N}(\bm{\mu_{Y}},\bm{\Sigma_{Y}})$ (24)
@@
$\begin{split}&\bm{{\mu}_{Z}}=\mathbb{E}[\bm{Z}]=(\mathbb{E}[Z_{1}],\mathbb{E}[Z_{2}],...,\mathbb{E}[Z_{k}]])^{T}\\ &\bm{{\mu}_{Y}}=\mathbb{E}[\bm{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\bm{\Sigma_{Z}}=\mathbb{E}[(\bm{Z}-\bm{{\mu}_{Z}})(\bm{Z}-\bm{{\mu}_{Z}})^{T}]\\ &\bm{\Sigma_{Y}}=\mathbb{E}[(\bm{Y}-\bm{{\mu}_{Y}})(\bm{Y}-\bm{{\mu}_{Y}})^{T}]\end{split}$ (25)
@@
Then the conditional distribution of $\bm{Y}$ given $\bm{Z}$ is also normally distributed Mood (1950) Johnson et al. (1995), which can be formulated as:
$\bm{Y}|\bm{Z}\sim\mathcal{N}(\mathbb{E}({\bm{Y}|\bm{Z}={Z}}),var(\bm{Y}|\bm{Z}={Z}))$ (26)
where $\mathbb{E}({\bm{Y}|\bm{Z}=Z})$ is the mean of the label set $\bm{Y}$ given a sample $\bm{Z}=Z$ from the embeddings, and $var(\bm{Y}|\bm{Z}=Z)$ is the variance of $\bm{Y}$ given a sample from the embeddings. The conditional mean $\mathbb{E}[(\bm{Y}|\bm{Z}=Z)]$ and conditional variance $var(\bm{Y}|\bm{Z}=Z)$ can be calculated as:
$\bm{\mu}_{\bm{Y}|\bm{Z}=Z}=\mathbb{E}[(\bm{Y}|\bm{Z}=Z)]=\bm{{\mu}_{Y}}+\bm{\Sigma_{YZ}}\bm{{\Sigma}^{-1}_{Z}}(Z-\bm{\mu_{Z}})$ (27)
$\bm{\Sigma}_{{\bm{Y|Z}}=Z}=var(\bm{Y}|\bm{Z}=Z)=\bm{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}$ (28)
where $\bm{\Sigma_{YZ}}$ and $\bm{\Sigma_{ZY}}$ are the cross-covariance matrices between $\bm{Y}$ and $\bm{Z}$, and between $\bm{Z}$ and $\bm{Y}$, respectively, and $\bm{\Sigma^{-1}_{Z}}$ denotes the inverse of the covariance matrix of $\bm{Z}$.